jigglypuff314 (jigglypuff314):

If \(f\) is a function such that \(f'(x) = -f(x)\), then \( ∫x f(x)~dx\) = a) x f(x) + f(x) + C b) -xf(x) - f(x) + C c) (x^2 /2) f(x) + C d) -(x^2 /2) f(x) + C e) - (x^2 /2) f(x) - (x^3 /6) f(x) + C symbolic notations makes my brain not think straight... it's probably really easy, but I just don't see it >,< help? :3

OpenStudy (tkhunny):

Integration by parts. \(\int xf(x)\;dx = \int f(x)\;d\left(½x^2\right) = ½x^{2}f(x) - \int ½x^{2}f'(x)\;dx\) Does that get us anywhere?

jigglypuff314 (jigglypuff314):

so u = f(x) v = (1/2)x^2 du = -f'(x) dv = x dx f(x) (1/2)(x^2) + (1/2) ∫ x^2 f'(x) dx (x^2 /2)f(x) + (1/2) ∫x^2 f'(x) dx it gets us closer to Options C, D, or E ... so then should I integrate by parts again? :/ @tkhunny

jigglypuff314 (jigglypuff314):

or would it keep on going forever... ? .-. whut

OpenStudy (sidsiddhartha):

im getting B

jigglypuff314 (jigglypuff314):

mmm could you explain how please? :3

OpenStudy (sidsiddhartha):

yeah

OpenStudy (sidsiddhartha):

\[f^{'}(x)=-f(x)\] \[\int\limits_{}^{}f^{'}(x)=-\int\limits_{}^{}f(x) \rightarrow f(x)=-\int\limits_{}^{}f(x)\] can we write it?

jigglypuff314 (jigglypuff314):

oh... that works... but how do I use that in my question?

OpenStudy (anonymous):

You could solve the differential equation \[f'(x)+f(x)=0\] for \(f(x)\) and then find \(\int xf(x)~dx\)

jigglypuff314 (jigglypuff314):

f(x) = f'(x) ? is that what you mean by "solve" ? so ∫x(f'(x)) dx then by parts? :3

OpenStudy (sidsiddhartha):

now \[\int\limits_{}^{}xf(x)dx=\int\limits_{}^{}-xf^{'}(x)dx\] \[=-[x \int\limits_{}^{}f^{'}(x)-\int\limits_{}^{}f(x)]=-[xf(x)+f(x)+c]\]

OpenStudy (anonymous):

My suggestion is probably beyond the scope of the question. Just another way to solve this. Is this for Calc II or some equivalent?

jigglypuff314 (jigglypuff314):

yes, Calc 2 / APCalcBC

OpenStudy (sidsiddhartha):

do u got my solution @jigglypuff314

OpenStudy (anonymous):

@sidsiddhartha, how is \(\int f(x)~dx=f(x)\)?

OpenStudy (anonymous):

Wait nevermind, I see what you did

OpenStudy (sidsiddhartha):

it is given f'(x)=-f(x) i have integrated both sides

OpenStudy (anonymous):

Yeah I noticed just now, thanks

OpenStudy (sidsiddhartha):

after that just used by parts then just substituted \[-\int\limits_{}^{}f(x)=f(x)\]

OpenStudy (anonymous):

Anyway @jigglypuff314 my suggestion involves the differential equation \[f'(x)+f(x)=0\] which gives the characterstic equation \[r+1=0~~\Rightarrow~~r=-1\] So the general solution is \[f(x)=Ce^{-x}\] And you can see that \[\begin{cases}f'(x)=-Ce^{-x}\\ f(x)=Ce^{-x}\end{cases}~~\Rightarrow~~f'(x)=Ce^{-x}=-Ce^{-x}=-f(x)\]

OpenStudy (anonymous):

Then if you integrated, \[C\int xe^{-x}~dx=-Cxe^{-x}+\int e^{-x}~dx=-Cxe^{-x}-Ce^{-x}+C^*=-xf(x)-f(x)+C^*\]

jigglypuff314 (jigglypuff314):

.-. um yeah I try sids' way O-O thanks @SithsAndGiggles xD thanks @sidsiddhartha I think I understand now, thank you all! <3